A thick walled steel cylinder used for storing compressed air is fitted with a safety valve which lifts at a pressure of 1.0x106 N/m2. It contains air at 17oC and 0.8x106 N/m2. At what temperature will the valve lift?
1
Expert's answer
2018-06-04T08:28:08-0400
Temperature and pressure of ideal gas of constant mass and volume are related by Gay-Lussac's law - T_1/T_2 =p_1/p_2 , where T_1,p_1 – are temperature and pressure of “cold” gas, and T_2 〖,p〗_2 - are temperature and pressure of “hot” gas. So T_2=(T_1 p_2)/p_1 =(290×1×〖10〗^6)/(0.8×〖10〗^6 )=290/0.8=362.5°K=89.5℃